OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b^2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0 Numerical Answers Expected!

OpenStudy (anonymous):

b^2 - 4ac, if the result is a square number it can be solved by factoring.

OpenStudy (anonymous):

can u help me please

OpenStudy (anonymous):

oh sorry ax^2 + bx + c = 0 2x^2 + 7x + 3 = 0 identify a, b and c then substitute in b^2 − 4ac, see if it's a square number

OpenStudy (anonymous):

bleh.. i need help

OpenStudy (anonymous):

D= ok a = 2 b = 7 c = 3 b^2 − 4ac = 7^2 - 4*2*3 = ?

OpenStudy (anonymous):

7^2 - 4(2)(3)

OpenStudy (anonymous):

so 14 - 24?

OpenStudy (anonymous):

7^2 isn't 14

OpenStudy (tyteen4a03):

The quadratic equation is \[x = \frac{ -b \pm \sqrt { b^2 - 4ac } }{ 2a }\] In the formula, if the \[b^2 - 4ac\] part is not a square number, we will not be able to factorize this equation without involving irrational numbers. Hence, \[b^2 - 4ac\] is the answer. In the second part, we have \[2x^2 + 7x + 3 = 0\] Substitute everything, you get \[2(7)^2 - 4(2)(3)\] Do the math and find out if it's a square number. --- Side note, the \[b^2 - 4ac\] expression is actually a discriminant to quadratic equations of this form \[ax^2 + bx + c = 0\]. It can be used to determine if the equation has 2 real roots, 1repeated root, or no real roots at all. If \[b^2 - 4ac > 0\] It means the equation has 2 real roots, If \[b^2 - 4ac = 0\] It means the equation has 1 repeated root. (Examples include (x-5)(x-5) = 0) If \[b^2 - 4ac < 0\] it means the equation has no real roots (it does not touch the x axis at all)

OpenStudy (tyteen4a03):

(Oops, I meant \[7^2 - 4(2)(3)\], not \[2(7^2) - 4(2)(3)\])

OpenStudy (anonymous):

so 49

OpenStudy (anonymous):

-24

OpenStudy (anonymous):

25 ?

OpenStudy (anonymous):

is 25 a square number?

OpenStudy (anonymous):

square number = 1^2, 2^2, 3^2, 4^2, 5^2, ... n^2

OpenStudy (anonymous):

no

OpenStudy (tyteen4a03):

To determine if a number is a square number, simply square root it and see if the number is a rational number. In this case, \[\sqrt{25}\] yields 5. 5 is a rational number, so it is a rational number.

OpenStudy (anonymous):

so the answer is 25

OpenStudy (anonymous):

(╯° Д° )╯︵ ┻━┻ןoן

OpenStudy (tyteen4a03):

(err, I mean that 25 is a square number)

OpenStudy (tyteen4a03):

The answer is not 25. It simply tells you that this equation can be factored. If you have trouble factorizing equations, I think you really need to check with your math teacher and revise factorization.

OpenStudy (anonymous):

tyteen the question didn't ask to solve it though. so 25 is what it asks for.

OpenStudy (anonymous):

so 25 is the answer :)

OpenStudy (anonymous):

(╯° Д° )╯︵ ɐɔɔǝqǝɹ jk